# Black hole surrounded by a magnetic vortex in $f(R)$ gravity

### v1

@@ III.2.2 On the Bekenstein-Hawking temperature
$-\frac{\partial^{2}\phi}{\partial t^{2}}+(r-r_{0})^{2}\frac{\partial^{2}\phi}{\partial r^{2}}+\frac{1}{2}\frac{\partial}{\partial r}(r-r_{0})^{2}\frac{\partial\phi}{\partial r}-m^{2}(r-r_{0})\phi=0.$ (36)
Assuming that the particle creation process in the BH is semiclassical in the BH background, the WKB approximation Sakurai informs us that the ansatz for the field $\phi$ Srinivasan , Angheben , Kerner1 , Mitra , Akhmedov is
@@
$\Theta(t,r)=-\omega t+W(r)$ (39)
In this scenario, $\omega$ is a constant representing the energy of the emitted radiation Lima .
@@
[…]
As a result, the probability of a particle escaping the BH via the tunneling process is
$\Gamma\sim\text{exp}(-2im\Theta)=\text{exp}(-4\pi\omega).$ (42)
Recalling that the tunneling probability (42) is related to the Boltzmann factor $\text{exp}(-\omega/T)$ Pathria , it follows that the Hawking temperature will be
@@
III.3 The topological vortex
We will focus on the study of magnetic vortices. To accomplish this, let us analyze the equations of motion within the context of modified $f(R)$ gravity (33), taking into account the quantum correction to the metric (13). In this framework, one notes that the equations of motion derived through the variation of the action concerning the scalar field, the gauge field, and the metric are, respectively,
@@ III.3 The topological vortex
Note that the set of linearly independent equations [(44)-(46)] is also invariant concerning the parameter $\alpha$. This behavior arises because cosmological magnetic vortices are classical objects not influenced by small quantum fluctuations originating from the quantum correction to the metric. Conversely, these vortices are significantly affected by the BH radius. That occurs because, as the event horizon expands, the vortex matter begins to collapse into the BH, causing increasing distortions in the magnetic vortex. We expose a representation of the collapse of the magnetic vortex matter into the BH in Fig. 1.
@@
[…]
Using the numerical interpolation approach Scarb , EBurden , let us investigate the numerical vortex solutions, i.e., the solutions for the field variables $g(r)$ and $a(r)$. We display the numerical solutions for the field variables $g(r)$ and $a(r)$, respectively, in Figs. 3(a) and 3(b).
(a) Field variable solutions of the matter sector.
(b) Field variable solutions of the gauge sector.
@@
Finally, evaluating the magnetic field [^10] One defines the intensity of the magnetic field of the vortex as $B=-F_{21}\equiv\frac{a^{\prime}(r)}{er}$ Weinberg . and energy density [^11] The energy density of the BH-vortex system is the $00$-component of the stress-energy tensor [$00$-component of Eq. (23)] Vachaspati , Rajaraman , Manton . of the BH-vortex system, we noted the existence of various magnetic oscillations near the event horizon of the BH, with energy density localized in the vortex. Naturally, the magnetic vortices formed in f(R) gravity are of the ring-like structures. We exposed the ring-like vortices in Figs. 4[(a)-(d)] and 5[(a)-(d)].
(a) Magnetic field intensity versus the vortex radius.
(b) Planar behavior for $r_{0}=0.1905$.
(c) Planar behavior for $\,r_{0}=0.2540$.
(d) Planar behavior for $\,r_{0}=0.2547$.
Figure 4: Magnetic field of the vortex vs. $r$.
(a) Energy density versus $r$.
(b) Planar behavior for $r_{0}=0.1905$.
(c) Planar behavior for $\,r_{0}=0.2540$.
(d) Planar behavior for $\,r_{0}=0.2547$.
Figure 5: Energy density of the BH-Vortex vs. $r$.
@@ IV Summary and conclusion
ACKNOWLEDGMENT
The authors are grateful to the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq). F. C. E. Lima and C. A. S. Almeida are supported, respectively, for grants No. 171048/2023-7 (CNPq/PDJ) and 309553/2021-0 (CNPq/PQ).
CONFLICTS OF INTEREST/COMPETING INTEREST
The authors declared that there is no conflict of interest in this manuscript.
DATA AVAILABILITY
No data was used for the research described in this article.

### v2

@@ II.0.1 On the Bekenstein–Hawking temperature
$-\frac{\partial^{2}\phi}{\partial t^{2}}+(r-r_{0})^{2}\frac{\partial^{2}\phi}{\partial r^{2}}+\frac{1}{2}\frac{\partial}{\partial r}(r-r_{0})^{2}\frac{\partial\phi}{\partial r}-m^{2}(r-r_{0})\phi=0$ .
Assuming that the particle creation process in the BH is semiclassical in the BH background, the WKB approximation informs us that the ansatz for the field $\phi$ Srinivasan , Angheben , Kerner1 , Mitra , Akhmedov is
@@
$\Theta(t,r)=-\omega t+W(r),$ (27)
where $\omega$ is a constant representing the energy of the emitted radiation.
@@
$W(r)=2\pi i\omega+\text{(real contribution)},$ (29)
As a result, the probability of a particle escaping the BH via the tunneling process is $\Gamma\sim\text{exp}(-2im\Theta)=\text{exp}(-4\pi\omega).$ Recalling that the tunneling probability is related to the Boltzmann factor $\text{exp}(-\omega/T)$ Pathria, it follows that the Hawking temperature will be
@@
III The topological vortex
We will focus on the study of magnetic vortices. To accomplish this, let us analyze the equations of motion within the context of modified $f(R)$ gravity (22). In this framework, one notes that the equations of motion derived through the variation of the action concerning the scalar field, the gauge field, and the metric are, respectively,
@@ III The topological vortex
Note that the set of linearly independent equations [(31)-(33)] is also invariant concerning the parameter $\alpha$. This behavior arises because cosmological magnetic vortices are classical objects not influenced by small quantum fluctuations originating from the quantum correction to the metric. Conversely, these vortices are significantly affected by the BH radius. That occurs because, as the event horizon expands, the vortex matter begins to collapse into the BH, causing increasing distortions in the magnetic vortex. We expose a representation of the collapse of the magnetic vortex matter into the BH.
@@
Here, $\nu$ is the VEV. For our analysis, we adopt, for simplicity, $\lambda=\beta=\nu=n=1$  [^9] Once $\nu$ describes the asymptotic value of the complex scalar field, i.e., $g(\infty)=\nu$, its modification only affects the value of $g(\infty)$, as announced in Eq. (34). Meanwhile, the parameter $\lambda$ becomes the topological structures more (or less) localized. Similarly, the parameter $\beta$ modifies the asymptotic value of the gauge field, i.e., $a(\infty)$ [see Eq. (7)]. Thus, for convenience, when studying topological vortex configurations, it is common to assume these parameters to be unitary, i.e., $\nu=\lambda=\beta=1$, once they do not produce any physical repercussions. For further details, see Ref. FLima3 .. Additionally, to examine the numerical solutions of the system of equations [(31)-(33)], we will use the numerical interpolation method, which allows us to estimate the solutions for the differential equations [Eqs. (31)-(33)]. That will provide us with the solutions for the field variables $g(r)$ and $a(r)$ that satisfy the Eqs. [(31)-(33)]. Generally speaking, we apply the numerical interpolation method and discretize the function’s domain. For instance, by considering the differential equations [(31)-(33)] defined in the range of the radial variable $r$, this range is subdivided into discrete steps. Thus, we discretize the position of the system by dividing the position range into two hundred thousand discrete points. Therefore, the solutions of the equations are investigated within the range $[0,200]$, using the discretization $r_{1},\,r_{2},\,\dots,\,r_{200000}$, where $r_{1}=0$ and $r_{200000}=200\,$ [^10] For further details on the numerical method, see Refs. Scarb , EBurden ..
Using the numerical interpolation approach Scarb , EBurden , let us investigate the numerical vortex solutions, i.e., the solutions for the field variables $g(r)$ and $a(r)$. We display the numerical solutions for the field variables $g(r)$ and $a(r)$, respectively, in Figs. LABEL:fig1(a) and LABEL:fig1(b).
(a) Field variable of the matter sector.
(b) Field variable of the gauge sector.
@@
Finally, evaluating the magnetic field [^11] One defines the intensity of the magnetic field of the vortex as $B=-F_{21}\equiv\frac{a^{\prime}(r)}{er}$ Weinberg . and energy density [^12] The energy density of the BH-vortex system is the $00$-component of the stress-energy tensor [$00$-component of Eq. (11)] Vachaspati , Rajaraman , Manton . of the BH-vortex system, we noted the existence of various magnetic oscillations near the event horizon of the BH, with energy density localized in the vortex. Naturally, the magnetic vortices formed in f(R) gravity are of the ring-like structures. We exposed the ring-like vortices in Figs. 3(a) and 3(b).
(a) Magnetic field vs. $r$.
(b) Energy density vs. $r$.
Figure 3: The numerical vortex solutions.
@@ IV Summary and conclusion
V ACKNOWLEDGMENT
The authors are grateful to the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq). F. C. E. Lima is supported, respectively, for grants No. 171048/2023-7 (CNPq/PDJ) and 309553/2021-0 (FAPESP/PQ).